Test the claim that the mean GPA for student athletes is significantly different than 3.3 at the 0.01 significance level.

Based on a sample of 75 people, the sample mean GPA was 3.28 with a standard deviation of 0.06

The test statistic is:  (to 3 decimals)

The p-value is:  (to 3 decimals)

Based on this we:

  • Reject the null hypothesis
  • Fail to reject the null hypothesis
Answer #1

Test statistic

3.28 – 3.3 0.06/ 75

t = -2.89

Degrees of freedom = 75 - 1 = 74

So,

p - Value = 2*P(t(74) < -2.89) = 0.005

We reject the null hypothesis.